OpenStudy (anonymous):

PLEASE HELP Let f(x)=-4x+7 and g(x)=10x-6. Find f(g(x)) Answer choices: A. -40x+24 B. -40x+31 C. -40x+64 D. -40x+70

OpenStudy (anonymous):

B

OpenStudy (anonymous):

f(g(x)) = 4.g(x) + 7 = -4(10x-6) + 7 = -40x+24+7=-40x+31 f(g(x)) = -40x + 31

OpenStudy (anonymous):

Thanks guys